# STENCIL: Submodular Mutual Information Based Weak Supervision for Cold-Start Active Learning

### v1

@@ 5 Conclusion
In this work, we present STENCIL, which effectively utilizes SMI and prior knowledge of the data environment in the form of a small exemplar set of rare-class instances to manage class-imbalanced cold-start active learning scenarios. We demonstrate that STENCIL is able to make improvements of $10\%-24\%$ overall accuracy and $17\%-40\%$ rare-class F-1 score over common active learning strategies within a single round of selection. As an exciting avenue of future work, the application of STENCIL’s SMI functionality to other text-based tasks beyond classification – and the formation of query exemplars in these tasks – would greatly expand the contributions of this work.
@@ Appendix A Chosen Query Exemplars within Experiments
In this section, we provide the query exemplars used in Section 4. The query phrases are created using domain knowledge, examples from prior work [16], and prompting ChatGPT [35] with the following: "Give me 10 example phrases, not sentences, found in each of 1. Spam youtube comments 2. Spam SMS messages 3. Tweets with positive sentiment." Notably, we did not examine the instances of YouTube Spam [29], SMS Spam [30], or Twitter Sentiment [31] when curating these query sets.
@@ A.2 SMS Spam
  • "Click this link to claim your cash prize"
  • "Confirm your account details to continue using our service"
  • "Congratulations! You’ve won a free trial"
  • "Earn money quickly"
  • "Exclusive discount code inside"
  • "Free gift awaiting you"
  • "Get paid to work from home"
  • "Important security notice"
  • "Problem with your payment method"
  • "Special limited-time offer"
  • "Suspicious activity detected"
  • "To stop receiving these messages, click here"
  • "Unclaimed money in your name"
  • "Urgent action required to receive your package"
  • "Verify your email and password immediately"
  • "Warning: Your account will be deactivated"
  • "You are eligible for a refund"
  • "You have an unpaid bill"
  • "You’re our lucky shopper today"
  • "You’ve been selected for an exclusive offer"
  • "Your account has been temporarily locked"
  • "Your order is ready for pickup"
  • "Your subscription is about to expire"
  • "Your trial period is ending"
  • "Your warranty is expired"
@@ A.3 Twitter Sentiment
  • "Absolutely love"
  • "Amazing job"
  • "Can’t wait for"
  • "Extremely happy to"
  • "Feeling blessed"
  • "Feeling inspired by"
  • "Feeling optimistic about"
  • "Feeling very proud of"
  • "Had a great time"
  • "Highly recommend"
  • "Incredible experience"
  • "Overwhelmed with happiness"
  • "Really excited about"
  • "So grateful for"
  • "Such a beautiful"
  • "Thank you so much"
  • "Totally loving the"
  • "Truly amazing"
  • "Very successful"
  • "Wonderful day"
@@ Appendix B Dataset Details
YouTube Spam Classification [29]: A collection of real comments on five of the ten most viewed YouTube videos at the time of collection. It consists of 1005 comments marked spam and 951 comments marked ham (non-spam). This dataset is publicly available in the UCI Machine Learning Repository [^1] https://archive.ics.uci.edu/dataset/380/youtube+spam+collection. After creating a balanced test set and making the spam class as the rare class in the training data, we induce a class imbalance of approximately $1:10$.
SMS Spam Classification [30]: A dataset of SMS messages curated for binary spam classification task. It combines 425 SMS spam messages from the Grumbletext website, a subset of 3,375 ham messages from the NUS SMS Corpus, 450 ham messages from a PhD Thesis, and 1,002 ham and 322 spam messages from the SMS Spam Corpus v.0.1 Big. This dataset is publicly accessible at the UCI ML Repository [^2] https://archive.ics.uci.edu/dataset/228/sms+spam+collection. After creating a balanced test set, the imbalance factor is left as approximately $1:18$.
Twitter Sentiment [31]: A sentiment classification dataset of comments from Twitter (tweets) about airline services. It includes 12,864 tweets labeled with positive, neutral, or negative sentiments based on analysis using an ensemble of 6 classifiers. For the scope of this study, only entries with positive and negative labels have been retained. The dataset is made available in [16]. After creating a balanced test set, the imbalance factor is left as approximately $1:6$ (with positive sentiments deemed rare).
@@ C.1 Submodular Mutual Information
As mentioned in Section 2, SMI can be instantiated utilizing various submodular functions, each of which models different properties. Facility Location ($F(A)=\sum_{i\in\mathcal{V}}\max_{j\in A}s_{ij}$) captures representation information between a subset and its ground set and is used to instantiate the FLVMI and FLQMI variants in Table 1. As derived in [20], FLQMI differs from FLVMI by modeling only the cross similarities between $A$ and $Q$. Log Determinant ($F(A)=\log\det S_{A}$) captures diversity information in $A$, which is reflected by the determinant of $S_{A}$; hence, LOGDETMI captures diversity information between $A$ and $Q$. Graph Cut ($F(A)=\sum_{i\in\mathcal{V}}\sum_{j\in A}s_{ij}-\lambda\sum_{i,j\in A}s_{ij}$) also focuses on representation of $\mathcal{V}$ with $A$ and, when ingested by SMI’s definition, gives GCMI, which focuses entirely on how relevant $A$ is to the queries in $Q$. Notably, all instantiations utilize similarity values between instances. In our experiments, we represent each instance via the average of the GloVe [25] embeddings for each token in the sequence. Afterwards, we compute cosine similarity to derive each $s_{ij}$ value.
@@ C.2 BADGE Sampling
The BADGE strategy [33] is a common choice for batch active learning where diverse sets of uncertain instances are selected, which tend to be informative instances. BADGE achieves this by embedding each unlabeled instance using hypothesized loss gradients (using the most confidently predicted class). By picking a diverse span of points with large magnitude gradients, one obtains a diverse set of instances that are likely to bring large updates to the loss (and are thus uncertain). In general, the following steps are taken:
@@
  • Using these gradients as a form of embedding for each unlabeled point, run k-means++ initialization [32] on this embedding set, retrieving $B$ centers. Each center is a point from the unlabeled set, and $B$ represents the active learning budget.
@@ C.3 Basic Uncertainty Sampling
A common and long-standing choice for active learning selection is that of simple uncertainty sampling measures as discussed in [5]. Namely, one can quantify uncertainty utilizing the predicted class probabilities and select the top $k$ most uncertain points. Typically, there are three common choices for quantifying uncertainty: Entropy ($H(x)=-\sum_{i}p(x)_{i}\log p(x)_{i}$), Least Confidence ($C(x)=\max_{i}p(x)_{i}$), and Margin ($M(x)=p(x)_{\sigma_{1}}-p(x)_{\sigma_{2}}$, where $\sigma_{1}$ and $\sigma_{2}$ denote the most probable and second-most probable class, respectively). Entropy sampling chooses the top $B$ $H(x)$ values. Least confidence sampling and margin sampling choose the $B$ smallest $C(x)$ and $M(x)$ values, respectively. The uncertain points are then labeled, which tends to provide more crucial information to the downstream model.
@@ C.4 RegEx Sampling
Appendix D Standard Deviations
In this section, we provide the standard deviation in metrics across the 10 trials whose averages are reported in Table 2 and Table 3. We present the standard deviations in Table 5 and Table 6.
Appendix E Reproducibility and Licenses
Dataset FLQMI FLVMI LDMI GCMI RegEx Rand Badge Ent L Conf Marg KMean
YouTube
70.9 $\pm$ 6.7
67.4 $\pm$ 3.4
77.8 $\pm$ 3.2
58.5 $\pm$ 5.7
66.6 $\pm$ 4.7
51.5 $\pm$ 5.5
53.3 $\pm$ 4.1
52.8 $\pm$ 8
52.1 $\pm$ 4.8
51.2 $\pm$ 3.1
56.3 $\pm$ 2.8
SMS
58.9 $\pm$ 6.9
66.6 $\pm$ 9.4
72.3 $\pm$ 4.9
86.9 $\pm$ 2.3
62.1 $\pm$ 12
55.8 $\pm$ 7.8
60.6 $\pm$ 6.9
68.2 $\pm$ 15
54.6 $\pm$ 9.4
60.2 $\pm$ 15
53.5 $\pm$ 3.6
Tweet
68.4 $\pm$ 3.2
66.8 $\pm$ 3.1
72.1 $\pm$ 1.6
66.2 $\pm$ 3.3
56.2 $\pm$ 4.0
59.1 $\pm$ 3.0
56.1 $\pm$ 4.4
53.2 $\pm$ 4.0
52.4 $\pm$ 5.3
53.4 $\pm$ 4.1
62.4 $\pm$ 2.8
YouTube
73.4 $\pm$ 5.8
55.1 $\pm$ 7.5
77.0 $\pm$ 5.9
61.1 $\pm$ 4.2
59.9 $\pm$ 7.5
10.0 $\pm$ 16
16.6 $\pm$ 13
12.8 $\pm$ 22
12.1 $\pm$ 15
9.7 $\pm$ 10.9
26.8 $\pm$ 7.6
SMS
27.9 $\pm$ 20
46.9 $\pm$ 22
61.1 $\pm$ 10
86.9 $\pm$ 2.2
33.4 $\pm$ 30
17.5 $\pm$ 21
32.8 $\pm$ 19
46.9 $\pm$ 34
12.7 $\pm$ 24
25.3 $\pm$ 36
11.8 $\pm$ 11.4
Tweet
72.4 $\pm$ 1.4
55.0 $\pm$ 7.3
67.2 $\pm$ 3.0
66.7 $\pm$ 4.4
24.3 $\pm$ 13
33.2 $\pm$ 8.9
24.0 $\pm$ 15
14.2 $\pm$ 14
11.2 $\pm$ 18
16.1 $\pm$ 14
43.3 $\pm$ 7.2
Dataset Sz-20% Sz-40% Sz-60% Sz-80% Sz-100%
YouTube
68.8 $\pm$ 8.8
70.8 $\pm$ 6.0
74.6 $\pm$ 4.0
74.7 $\pm$ 2.5
77.2 $\pm$ 3.3
SMS
63.9 $\pm$ 7.3
68.8 $\pm$ 4.2
70.9 $\pm$ 8.3
70.3 $\pm$ 8.4
73.2 $\pm$ 4.5
Tweet
69.5 $\pm$ 3.7
70.5 $\pm$ 2.9
71.6 $\pm$ 1.9
71.9 $\pm$ 1.6
72.3 $\pm$ 1.6
YouTube
58.4 $\pm$ 21.4
63.6 $\pm$ 12.6
71.9 $\pm$ 6.2
71.8 $\pm$ 5.9
76.9 $\pm$ 6.5
SMS
41.5 $\pm$ 18.8
54.1 $\pm$ 9.5
57.2 $\pm$ 17.8
55.9 $\pm$ 20.2
63.1 $\pm$ 8.9
Tweet
62.0 $\pm$ 8.5
63.6 $\pm$ 6.0
67.0 $\pm$ 3.7
67.1 $\pm$ 3.7
67.5 $\pm$ 2.7
  • DISTIL [6]: MIT License
  • Submodlib [36]: MIT License
  • PyTorch [37]: Modified BSD License
  • NLTK [38]: Apache License Version 2.0
  • Transformers [39]: Apache License 2.0
  • YouTube Spam [29]: CC Attribution 4.0 International
  • SMS Spam [30]: CC Attribution 4.0 International
  • Twitter Sentiment [31]: Creative Commons CC-BY-NC-ND
Appendix F Limitations and Societal Impacts

### v2

@@ 5 Conclusion
In this work, we present STENCIL, which effectively utilizes SMI and prior knowledge of the data environment in the form of a small exemplar set of rare-class instances to manage class-imbalanced cold-start active learning scenarios. We demonstrate that STENCIL is able to make improvements of $10\%-18\%$ overall accuracy and $17\%-40\%$ rare-class F-1 score over common active learning strategies within a single round of selection. As an exciting avenue of future work, the application of STENCIL’s SMI functionality to other text-based tasks beyond classification – and the formation of query exemplars in these tasks – would greatly expand the contributions of this work.
Reproducibility and Licenses
  • DISTIL (Beck et al., 2021): MIT License
  • Submodlib (Kaushal et al., 2022): MIT License
  • PyTorch (Paszke et al., 2019): Modified BSD License
  • NLTK (Bird et al., 2009): Apache License Version 2.0
  • Transformers (Wolf et al., 2020): Apache License 2.0
  • YouTube Spam (Alberto et al., 2015): CC Attribution 4.0 International
  • SMS Spam (Almeida et al., 2011): CC Attribution 4.0 International
  • Twitter Sentiment (Wan & Gao, 2015): Creative Commons CC-BY-NC-ND
@@ Appendix A Chosen Query Exemplars within Experiments
In this section, we provide the query exemplars used in Section 4. The query phrases are created using domain knowledge, examples from prior work (Maheshwari et al., 2022), and prompting ChatGPT (OpenAI, 2023) with the following: ”Give me 10 example phrases, not sentences, found in each of 1. Spam youtube comments 2. Spam SMS messages 3. Tweets with positive sentiment.” Notably, we did not examine the instances of YouTube Spam (Alberto et al., 2015), SMS Spam (Almeida et al., 2011), or Twitter Sentiment (Wan & Gao, 2015) when curating these query sets.
@@ A.2 SMS Spam
  • ”Click this link to claim your cash prize”
  • ”Confirm your account details to continue using our service”
  • ”Congratulations! You’ve won a free trial”
  • ”Earn money quickly”
  • ”Exclusive discount code inside”
  • ”Free gift awaiting you”
  • ”Get paid to work from home”
  • ”Important security notice”
  • ”Problem with your payment method”
  • ”Special limited-time offer”
  • ”Suspicious activity detected”
  • ”To stop receiving these messages, click here”
  • ”Unclaimed money in your name”
  • ”Urgent action required to receive your package”
  • ”Verify your email and password immediately”
  • ”Warning: Your account will be deactivated”
  • ”You are eligible for a refund”
  • ”You have an unpaid bill”
  • ”You’re our lucky shopper today”
  • ”You’ve been selected for an exclusive offer”
  • ”Your account has been temporarily locked”
  • ”Your order is ready for pickup”
  • ”Your subscription is about to expire”
  • ”Your trial period is ending”
  • ”Your warranty is expired”
@@ A.3 Twitter Sentiment
  • ”Absolutely love”
  • ”Amazing job”
  • ”Can’t wait for”
  • ”Extremely happy to”
  • ”Feeling blessed”
  • ”Feeling inspired by”
  • ”Feeling optimistic about”
  • ”Feeling very proud of”
  • ”Had a great time”
  • ”Highly recommend”
  • ”Incredible experience”
  • ”Overwhelmed with happiness”
  • ”Really excited about”
  • ”So grateful for”
  • ”Such a beautiful”
  • ”Thank you so much”
  • ”Totally loving the”
  • ”Truly amazing”
  • ”Very successful”
  • ”Wonderful day”
@@ Appendix B Dataset Details
YouTube Spam Classification (Alberto et al., 2015): A collection of real comments on five of the ten most viewed YouTube videos at the time of collection. It consists of 1005 comments marked spam and 951 comments marked ham (non-spam). This dataset is publicly available in the UCI Machine Learning Repository [^2] https://archive.ics.uci.edu/dataset/380/youtube+spam+collection. After creating a balanced test set and making the spam class as the rare class in the training data, we induce a class imbalance of approximately $1:10$.
SMS Spam Classification (Almeida et al., 2011): A dataset of SMS messages curated for binary spam classification task. It combines 425 SMS spam messages from the Grumbletext website, a subset of 3,375 ham messages from the NUS SMS Corpus, 450 ham messages from a PhD Thesis, and 1,002 ham and 322 spam messages from the SMS Spam Corpus v.0.1 Big. This dataset is publicly accessible at the UCI ML Repository [^3] https://archive.ics.uci.edu/dataset/228/sms+spam+collection. After creating a balanced test set, the imbalance factor is left as approximately $1:18$.
Twitter Sentiment (Wan & Gao, 2015): A sentiment classification dataset of comments from Twitter (tweets) about airline services. It includes 12,864 tweets labeled with positive, neutral, or negative sentiments based on analysis using an ensemble of 6 classifiers. For the scope of this study, only entries with positive and negative labels have been retained. The dataset is made available in (Maheshwari et al., 2022). After creating a balanced test set, the imbalance factor is left as approximately $1:6$ (with positive sentiments deemed rare).
@@ C.1 Submodular Mutual Information
As mentioned in Section 2, SMI can be instantiated utilizing various submodular functions, each of which models different properties. Facility Location ($F(A)=\sum_{i\in\mathcal{V}}\max_{j\in A}s_{ij}$) captures representation information between a subset and its ground set and is used to instantiate the FLVMI and FLQMI variants in Table 1. As derived in (Kothawade et al., 2022c), FLQMI differs from FLVMI by modeling only the cross similarities between $A$ and $Q$. Log Determinant ($F(A)=\log\det S_{A}$) captures diversity information in $A$, which is reflected by the determinant of $S_{A}$; hence, LOGDETMI captures diversity information between $A$ and $Q$. Graph Cut ($F(A)=\sum_{i\in\mathcal{V}}\sum_{j\in A}s_{ij}-\lambda\sum_{i,j\in A}s_{ij}$) also focuses on representation of $\mathcal{V}$ with $A$ and, when ingested by SMI’s definition, gives GCMI, which focuses entirely on how relevant $A$ is to the queries in $Q$. Notably, all instantiations utilize similarity values between instances. In our experiments, we represent each instance via the average of the GloVe (Pennington et al., 2014) embeddings for each token in the sequence. Afterwards, we compute cosine similarity to derive each $s_{ij}$ value.
@@ C.2 BADGE Sampling
The BADGE strategy (Ash et al., 2019) is a common choice for batch active learning where diverse sets of uncertain instances are selected, which tend to be informative instances. BADGE achieves this by embedding each unlabeled instance using hypothesized loss gradients (using the most confidently predicted class). By picking a diverse span of points with large magnitude gradients, one obtains a diverse set of instances that are likely to bring large updates to the loss (and are thus uncertain). In general, the following steps are taken:
@@
  • Using these gradients as a form of embedding for each unlabeled point, run k-means++ initialization (Arthur et al., 2007) on this embedding set, retrieving $B$ centers. Each center is a point from the unlabeled set, and $B$ represents the active learning budget.
@@ C.3 Basic Uncertainty Sampling
A common and long-standing choice for active learning selection is that of simple uncertainty sampling measures as discussed in (Settles, 2009). Namely, one can quantify uncertainty utilizing the predicted class probabilities and select the top $k$ most uncertain points. Typically, there are three common choices for quantifying uncertainty: Entropy ($H(x)=-\sum_{i}p(x)_{i}\log p(x)_{i}$), Least Confidence ($C(x)=\max_{i}p(x)_{i}$), and Margin ($M(x)=p(x)_{\sigma_{1}}-p(x)_{\sigma_{2}}$, where $\sigma_{1}$ and $\sigma_{2}$ denote the most probable and second-most probable class, respectively). Entropy sampling chooses the top $B$ $H(x)$ values. Least confidence sampling and margin sampling choose the $B$ smallest $C(x)$ and $M(x)$ values, respectively. The uncertain points are then labeled, which tends to provide more crucial information to the downstream model.
@@ C.4 RegEx Sampling
Appendix D Limitations and Societal Impacts